### v1

# Vision-driven Autonomous Flight of UAV Along River Using Deep Reinforcement Learning with Dynamic Expert Guidance

Abstract
Vision-driven autonomous flight and obstacle avoidance of Unmanned Aerial Vehicles (UAVs) along complex riverine environments for tasks like rescue and surveillance requires a robust control policy, which is yet difficult to obtain due to the shortage of trainable river environment simulators and reward sparsity in such environments. To easily verify the navigation controller performance for the river following task before real-world deployment, we developed a trainable photo-realistic dynamics-free riverine simulation environment using Unity. Successful river following trajectories in the environment are manually collected and Behavior Clone (BC) is used to train an Imitation Learning (IL) agent to mimic expert behavior and generate expert guidance. Finally, a framework is proposed to train a Deep Reinforcement Learning (DRL) agent using BC expert guidance and improve the expert policy online by sampling good demonstrations produced by the DRL to increase convergence rate and policy performance. This framework is able to solve the along-river autonomous navigation task and outperform baseline RL and IL methods. The code and trainable environments are available [^1] https://github.com/lijianwen1997/Deep-Reinforcement-Learning-with-Dynamic-Expert-Guidance.
I INTRODUCTION
Autonomous flight of UAVs over inland waterways has been investigated in applications like water sampling [1], inspection of irrigation system [2], and cooperative drone-boat navigation tasks [3, 4, 5]. However, very few works addressed the along-river autonomous navigation tasks of UAVs, which can be applied in rescue, surveillance, and inspection missions over inland waterways. Due to seasonal changes of river channels’ geometry and water level, manually planned GPS waypoints may not suit for accurate river following missions. Besides, when navigating in cluttered riverine environments, UAVs also need to avoid obstacles like trees and bridges, while being above the water.
To address the above challenges, a camera is used as the major perception source for the UAV to take advantage of information embedded in images and the recent development in image processing. To reduce the cost and safety issues for real-world experiments, we developed a series of riverine simulation environments using Unity [6] to verify the feasibility of controlling a UAV to fly along a river with only photo-realistic vision input. The impact of environmental forces on UAV is not considered in the simulation environment and the UAV action in the 3D space is discretized to alleviate the navigation challenge of operating without a map along rivers.
Using the above simplification, we decouple the optimization of high-level local waypoint-based control policies or navigation planners for UAVs from mid-level velocity control, which is more suitable for the river following task due to its larger acceptable operation space compared to autonomous driving tasks for cars.
We pre-train an Imitation Learning (IL) agent using human demonstrations to provide basic guidance, then combine Reinforcement Learning (RL) algorithm with an online IL algorithm that accumulates and learns from decent agent experiences, to boost the learning speed and final agent performance. This will overcome the challenge of obtaining a performant controller using only RL algorithms since the only positive reward is given when the UAV traverses any newly visited discretized river segment. This training framework is examined to outperform Proximal Policy Optimization (PPO), Behavior Clone (BC), and PPO with static BC guidance.
In summary, the main contributions of this paper are:
  • The creation of the trainable photo-realistic riverine simulation environment, which bridges the sim2real gap for vision-driven autonomous navigation of UAVs in the river following task.
  • The development of Deep Reinforcement Learning with a dynamic expert guidance framework that interactively trains Reinforcement Learning and Imitation Learning agents to reduce the training time and increase the overall performance.
II RELATED WORK
Autonomous river following in a relatively simple curved river segment has been tested using a water segmentation mask from HSV threshed images [7, 8]. This image processing technique is preferable in terms of its implementation simplicity but has vulnerabilities like unverified generalization ability to rivers under different geometrical and light conditions, as well as the high safety hazards and experimental costs.
Game engines as simulation environments have brought more advantages for vision-driven Reinforcement Learning (RL) training, due to their flexibility in customizing environmental parameters, their low cost as a testbed before real-world deployment, and most importantly, their photo-realism to reduce the sim2real gap in air [9, 10], land [11, 12] (indoor/outdoor), and even underwater applications [13].
To achieve a safer and more robust autonomous navigation controller in a diverse and adaptive riverine domain, researchers utilized Unreal Engine to train the agent for the river following task [14, 15]. Their proof of the superiority of the Variational Auto Encoder (VAE) [16] for image encoding, over Linear Regression method or end-to-end controller training, is adopted in our work. However, the reset criteria and the demonstrative trajectory collection approaches of their river simulation environment are manual-based, the agent action is limited to 2D planes, and the environment is not publicly available. To overcome these difficulties, our Unity-based river simulator can be automatically trained and tested for the river following task, supports 3D action space and collision detection, and is made public as an OpenAI gym [17] environment, thanks to Unity ML-Agents Toolkit [6].
Reinforcement learning(RL) and Imitation Learning(IL) are two powerful techniques with significant applications in the field of robotics. Robots can learn optimal behavior through trial and error through RL. Robots can learn to avoid obstacles [18], plan paths [19], and perform tasks like autonomous driving [20] through RL even in complicated, dynamic situations. On the other hand, IL [21] leverages expert demonstrations to train robots. By monitoring and imitating the actions of experts, robots can acquire complex skills rapidly. This is valuable in scenarios where human expertise is readily available, as it allows robots to learn efficiently without extensive programming [22].
However, both of these approaches have some limitations. RL frequently requires a large number of interactions with the environment to learn optimal policies, which can be time-consuming and expensive. Recently, Researchers have combined RL and IL to accelerate the training of RL using IL’s guidance. For example, [23] proposed to use the imitative expert policy to guide the learning of the DRL agent by regularizing the divergence between the DRL agent’s policy and the expert policy. [24] proposed a framework that uses a loss function that combines behavior cloning and 1-step Q-learning losses to enable the transition from BC to RL without performance degradation. However, after the DRL agent has developed a policy that is similar to the expert policy if the DRL agent tries to update the policy to improve the performance, it will be constrained by the expert policy or the value penalty thus the improvement is hindered.
On the other hand, as they have predominantly learned to imitate the demonstrations, robots trained through IL may not be able to adapt well to unexpected circumstances. DAgger [25] improves on BC by training on a dataset that better resembles the observations the expert policy is likely to encounter. However, the IL agent is also constrained by the expert. Thus to overcome these limitations, we propose an interactive framework called DRL with dynamic expert guidance that trains an RL agent using expert guidance and improves the expert policy online by sampling demonstrations produced by the DRL expert so that IL and RL can be improved jointly.
(a) Comparison of image captured in Wildcat Creek, Indiana, USA (left) and image from Unity river environment (right) that are alike in overall components layout and texture appearance.
(b) Orthographic images of the training (left) and testing (right) river simulation environments.
Figure 1: Real and simulation river environments.
III SIMULATION
III-A Photo-realistic River Environments
The riverine simulation environment is built in Unity using River Auto Material 2019  [^2] https://assetstore.unity.com/packages/tools/terrain/r-a-m-2019-river-auto-material-2019-145937 and Mountain Environment  [^3] https://assetstore.unity.com/packages/3d/vegetation/mountain-environment-dynamic-nature-191851. The river spline tool is provided by the former package to support fast, flexible and scriptable river building with various realistic water textures and shaders. The latter package provides a huge library of scanned forest assets like trees, branches, logs, and rocks, which helps to enrich the biomes of the riverine scene, as well as makes it more alike the real-world river scene components and appearance, as shown in Figure 0(a).
Aside from the realistic river and terrain trees, the riverine simulation environment features several additional aspects that not only mimic the general objects in real-world rivers, creeks, and streams, but also make the training of river following agents more difficult: 1) river channels with varied widths and depths, 2) river tributaries that diverge from the main water stream, 3) bridges over the river that can obstruct the navigable path of UAVs, and 4) sharp and slow turns. (Figure 2)
Several environment design choices are made for better photo-realism and more robust training. Instead of continuous velocity control, discrete position control for UAV movement is used, which eliminates the efforts of controller dynamics adaptation in real-world riverine environments with various wind speeds, water surface effects, etc. Thus we will name the UAV agent as the camera agent in the contents hereafter. The circular shape of rivers in our simulated environments is designed for better camera view coherence without seeing too many unrealistic un-modeled parts (sky, terrains, etc.) as the camera agent moves to any valid pose along the river. Besides, in the circular river environment, no matter where the camera agent is reset at the beginning of a training episode, the agent policy always has a chance to learn to traverse the whole river before ending this episode, making the collected trajectory lengths more uniform and independent of agent reset positions.
(a) Varying widths and depths
(b) Tributary
(c) Bridge
(d) Challenging U turns
Figure 2: Key features of the Unity river simulation environment.
TABLE I: Multi-discrete action values of the camera agent in the actual execution order, whose controlling logic is the same as remote controllers of common UAVs.
Action 0 1 2
Up-down Translation (m) 1 0 -1
Horizontal Rotation (deg) 10 0 -10
Forward-backward Translation (m) 1 0 -1
Leftward-rightward Translation (m) 0.5 0 -0.5
III-B Camera Agent
The simulation agent, created using Unity ML-Agents Toolkit [6], is a camera that looks forward with a fixed pitch angle (20 degrees in our experiments) with respect to the drone frame. Since no dynamics is included in this simulation, and all control actions are discrete, the drone camera itself is viewed as the agent. The camera agent is controlled by a four-dimensional multi-discrete action that maps from the command inputs of the common remote controller of UAVs: 1) translation up and down - throttle, 2) horizontal rotation counter-clockwise and clockwise - yaw rate, 3) translation forward and backward - pitch, 4) translation leftward and rightward - roll. Each action branch has three discrete values $\{0,1,2\}$, where $1$ means no action on that branch. The actions are applied in the drone coordinate frame (instead of the camera frame) with the right-hand rule, where the positive X axis points along the drone head and the positive Z axis points upward to the sky. The specific values of movement caused by actions are in Table I.
III-C Gym Environment
A river can be abstracted as a 3D Catmull-Rom spline [26] that is interpolated from several control points. The river surface is formed if we stretch this spline bi-directionally along some direction. This closed spline is lifted up to form a 3D bounding volume to limit the height of the camera agent, as well as its horizontal position. The tangents along the spline provide failure judgment when the camera agent’s yaw angle is far from either tangent angle of the nearest point on this spline with the projected agent position. This prohibits the turnaround motion of the agent since the built river is circular so that the agent can finish the whole-river navigation without any turnaround. Even if the agent does not trigger the above two ”done” criteria, lingering around or doing continuous meaningless movements with no reward improvement (over 50 steps in our experiments) will also trigger environment reset, and so does collision with bridges. If the environment ”done” is triggered during training, the camera agent will be reset at a random point inside the spline-based bounding volume, with a random yaw near either spline tangent of the nearest spline point. The above criteria for the river following task can be better visualized in Figure 5.
Figure 3: Demonstrative diagram of valid activity space and acceptable yaw range of camera agent in the river following task. $h_{1}=1m,h_{2}=15m,\alpha=150^{\circ}$ in our experiments. The drone picture is from  [^5] https://www.jaycar.com.au/r-c-fpv-drone-with-1080p-camera/p/GT4118.
The river central spline can be simplified as a sequence of connected line segments with nearly the same length. To have a decent baseline RL algorithm to compare, and to intuitively sense the learning capabilities of different learning algorithms, shaped reward is given to the camera agent if the projection of its positions on the spline before and after an action covers any un-visited line segments. The line segment where the camera agent’s current position projects onto is denoted as visited. If the camera agent has visited all the river splines, it will get the highest reward, then the environment will be reset. For collision detection, a mesh collider is used for bridges, and a box collider with a side length of $0.5m$ is used for the camera agent. The overall reward function for the autonomous river following task is given below.
$r(s,a)=\begin{cases}-1,&\text{agent out of spline bounding volume}\\ -1,&\text{agent yaw diverts too much}\\ -1,&\text{agent collides with bridge}\\ -1,&\text{agent no progress over certain steps}\\ r_{shaped}&\text{agent covers un-visited spline segment}\\ 0,&\text{otherwise}\end{cases}$ (1)
where $r_{shaped}$ is the percentage of newly-visited segments $n$ after the action $a$ over all segments number $N$, multiplied by a constant value, $10$ in this case.
$r_{shaped}=10*n/N\$ (2)
All negative reward conditions denote the reset of the environment (Equation 1). Besides, no final positive reward is given when the camera agent has visited all river spline segments since the amplified shaped reward, Equation 2, already gives informative guidance for the river following task. Due to the same reason, no environment reset is applied for backup movement or staying still of the agent.
IV METHODOLOGY
Figure 4: System architecture. A human expert collects good trajectory before the training and the transitions are represented by blue arrows. $a^{H}_{t}$ denotes the human expert’s action. The transitions during training are represented by black arrows. $a^{E}_{t}$ denotes the IL expert’s action.
The proposed method is shown in Figure 4 and algorithm 1 to solve the river following task. It requires a demonstration dataset from a human expert, an IL agent to provide guidance, and an RL agent that explores the environment, updates the dataset and learns a control policy.
To train an IL agent to learn from demonstration, a dataset $\mathcal{D}^{E}:\{\tau_{i}\}_{N}$ for $N$ successful trajectories is needed. Each trajectory is a sequence of state-action pairs $\tau={s_{1},a_{1},...s_{t},a_{t}}$. The goal of the expert policy $\pi^{E}$ is to maximize the log-likelihood over $\mathcal{D}^{E}$.
[…]
To keep improving the expert policy’s performance during the DRL training, trajectories with episodic rewards larger than the reward threshold $T_{reward}$ will be appended to $\mathcal{D}^{E}$. $T_{reward}$ is $4$ in our experiments. Once the dataset is updated, the expert policy is retrained using the newest $2000$ steps from the dataset until it can mimic at least $95\%$ behaviors from the demonstration.
Algorithm 1 DRL with Dynamic Expert Guidance
Expert collects dataset $\mathcal{D}^{E}$
Train the IL policy $\pi^{E}$
Initialize policy network $\phi$, value network $\psi$
repeat
     Observe state $s_{t}$
     Sample actions $a_{t}\sim\pi^{\phi}$ and $a_{t}^{E}\sim\pi^{E}$
     Execute the action $a_{t}$ in the environment
     Observe the next state $s_{t+1}$ and reward $r_{t}$
     Store transition $(s_{t},a_{t},s_{t+1},r_{t})$ in the agent buffer $\mathcal{D}^{R}$
     if time to update DRL then
         Sample a mini-batch $\mathcal{B}$
         Calculate KL divergence $D_{KL}$
         Calculate loss $\mathcal{L}$
         Update value network $\psi$
         Update policy network $\phi$
     end if
     if time to evaluation DRL then
         Sample trajectories using $\pi^{\phi}$
         if $\tau$ has a reward above $T_{reward}$  then
              Append the trajectory $\tau$ in $\mathcal{D}^{E}$
              Update the IL policy $\pi^{E}$
         end if
     end if
until Convergence or maximum training steps
The target of RL is to obtain a policy, defined as a distribution over actions conditioned on states $\pi(a_{t}|s_{t})$, to maximize the long-term discounted cumulative rewards
[…]
where $\tau$ is a trajectory and $p_{\pi}(\tau)$ is the distribution of the trajectory under policy $\pi$, and $T$ is the time horizon. Due to Proximal Policy Optimization(PPO)’s scalability and generalization ability, we use PPO [27] as a DRL algorithm to update the parameters of the policy network. In our framework, we use policy loss, value loss, and additional action loss to calculate the total loss and optimize the learning parameters.
The policy loss constrains the biggest possible improvement step on a policy using the data we currently have.
$\mathcal{L}_{policy}=\mathbb{E}\Bigl{[}\mathrm{min}{\Bigl{(}r_{t}(\theta)\hat{A_{t}},clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A_{t}}\Bigr{)}}\Bigr{]}$ (5)
where $r_{t}(\theta)=\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{\theta}_{old}(a_{t}|s_{t})}$ denotes the probability ratio between the current and old policy. $\hat{A_{t}}$ is the advantage. The value function network $V_{\psi}$ is updated by the following loss function:
$\mathcal{L}_{value}=\mathbb{E}\Bigl{[}\Bigl{(}R_{t}-V_{\phi}(s_{t})\Bigr{)}^{2}\Bigr{]}$ (6)
where $R_{t}$ represents the returned value. To encourage the learned policy to be close to the policy learned from the demonstration, we can set a penalty term in the loss function
$\mathcal{L}_{action}=D_{KL}(\pi^{\theta},\pi^{E})=\sum\pi^{\theta}(a_{t}|s_{t})\log\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{E}(a_{t}|s_{t})}$ (7)
where $D_{KL}$ is the KL divergence that measures the divergence between two distributions.
Combining the above loss functions for the DRK with dynamic expert guidance becomes
$\mathcal{L}=c_{policy}\mathcal{L}_{policy}+c_{value}\mathcal{L}_{value}+c_{action}\mathcal{L}_{action}$ (8)
V SIMULATION VERIFICATION
One training and one testing river simulation environment 0(b) are created to train learning algorithms and test their generalization abilities. Specifically, these maps differ in the overall river length, variation of river widths and depths, number of (sharp) turns, bridges, tributaries, and sand islands, whereas the lighting conditions and terrain textures are unified to compare the performance of vision-driven controllers in learning geometrical patterns for the river following task.
Observation images are resized to $128$ x $128$ for both encoding and RL training, and the encoded image is a vector of $1024$ float numbers. Variational Autoencoder [16] is obtained for each map by training 2000 randomly collected images from all maps up to the current training map’s difficulty level. This progressive update of the image encoder depends on the texture coherence across all maps, reduces the impact of the encoding module on RL training, and accounts for the downward compatibility of the trained agent. The encoded mean vector (without Gaussian noise) is used as the observation input to RL algorithms.
We ask an expert to execute the designed river following tasks in the simulator and collect actions and states as expert demonstration data. The UAV is randomly initialized in the valid activity space. The expert tries to correct the pose of the UAV and then controls the UAV to follow the river. Each trajectory is $50$ steps long. We collect $50$ trajectories in the medium environment as $\mathcal{D}^{E}$. The expert policy $\pi^{E}$ is trained by behavior clone using $\mathcal{D}^{E}$. After $50$ epochs of training, the expert policy $\pi^{E}$ has a success rate of $93.27\%$ to predict the expert action from the demonstration and is able to achieve a mean reward of $0.50$ in the training environment. The expert policy is used to generate reference actions during the training of RL agents.
Training is done on a desktop computer with a GeForce 2080 GPU and an Intel Core i7-8700 CPU. Code related to RL and IL is based on StableBaseline3 [28] and Imitation Learning Baseline Implementations [29].
VI RESULTS AND ANALYSIS
VI-A Training Results
The proposed methods together with baseline methods (BC, PPO, and PPO+StaticBC) are used to train a camera agent in the river following environment. All the methods have been trained for $14000$ steps for $2$ different seeds.
Figure5 shows the training results of different methods in the training environment. The training curve is generated by averaging the training results from different seeds and represented by a solid line of the mean reward with ±$\pm 1$ standard error. The episodic reward is calculated by averaging $100$ past episodes and updated every $1024$ steps. The blue horizon line represents the average of $100$ episodic rewards of the expert policy, which is obtained from 50 episodes of evaluation in $2$ different seeds.
PPO baseline slowly improves throughout the reinforcement learning phase but is still lower than the pre-trained BC baseline. The PPO+DaynamicBC initially returns rewards similar to PPO+StaticBC baseline. Both methods can get a reward at or above the BC after $25000$ steps. However, after the DRL agent’s performance surpasses the BC, PPO+DaynamicBC’s learning curve shows a higher increasing speed and finally reaches a mean episodic reward of $2.0$ at $105000$ steps, while the best mean episodic reward of PPO+StaticBC is less than $1.5$. PPO+StaticBC also shows multiple decrease trends during training, mainly because of the constraint of the static expert policy. PPO+DaynamicBC shows one decrease trend ($110000$ to $13000$) however it recovered to $1.7$ after $10000$ steps. On average, PPO+DaynamicBC’s score is $1.5$ higher than PPO, $0.6$ higher than the BC, and $0.4$ higher than the PPO+StaticBC.
Figure 5: Training results of different approaches in the river following environment. The green horizontal line represents the performance of the pre-trained BC.
VI-B Testing Results
Figure 6: Episode reward distributions of four algorithms on training and testing environments. The horizontal bar represents the mean, the vertical bar represents 25% and 75% percentiles, and the white dot is the median. Statistics from 50 episodes for each algorithm.
Although both algorithms PPO+StaticBC and PPO+DynamicBC can achieve high episodic reward in the training map, the dynamic one still achieves higher mean and median episodic rewards in both maps and leads the static one in terms of the maximum episodic reward by a large margin in the testing map (Figure 6). This verifies the superiority of PPO+DynamicBC algorithm since the IL (BC) model is evolving itself while guiding RL (PPO) model, which acts as a better regularizer compared to the PPO+StaticBC algorithm. PPO baseline and BC algorithms perform the worst in both training and testing maps, due to shortage of runtime expert guidance and limited generalization ability, respectively.
Figure 7: Trajectories of two algorithms PPO+StaticBC and PPO+DynamicBC starting from the same pose in the training map, in two experiments. Star symbols signify the agent position every 50 steps. Trajectory color gets darker as the agent moves further away from the starting position.
To qualitatively compare the navigation efficiency and failure modes of both dynamic and static algorithms, we plot their trajectories in the training map when starting at the same locations (Figure 7). The static algorithm agent can only navigate about $70$% of river length in both experiments, before it failed due to exceeding river spline volume (distracted by the tributary 0(b)) and bumping into the bridge, respectively. This is better visualized by the agent view near blue circles. The dynamic algorithm agent not only finishes whole-river navigation in both experiments but also with fewer steps, which justifies its robustness and efficiency. This can be visualized by the leading checkpoints (star in the figure) of the dynamic algorithm agent, meaning the agent moves a longer effective distance with fewer meaningless movements within the same 50 steps than the static agent. Besides, the dynamic algorithm handles better around river corners with smoother trajectory and larger safe margins to river banks. It also rejects tributary when seeing the main stream ahead by timely heading correction. Both cases can be viewed in figures near the red circles in Figure 7.
VII CONCLUSION
In this paper, we develop two riverine simulation environments to train a UAV agent to autonomously follow the river and avoid obstacles. A novel reinforcement learning method that utilizes the dynamic IL expert has been proposed. The performance of the proposed method has been compared with the baseline methods (PPO, BC, and PPO+StaticBC) in the riverine environments and achieves good qualitative and quantitative results. In the training environment, it achieves at least $0.4$ more rewards than all other methods on average during training and the highest best reward, and it is able to efficiently finish the river course without collision. It has good generalization ability as it also performs best in an unseen and difficult testing environment.
In the near future, tunable environmental light conditions like light directions and intensities can be enabled in the simulated riverine gym environment to increase policy robustness in the way of domain randomization and transfer the learned river following ability to a real UAV platform and test it in real-world riverine environment. On the other hand, we will systematically study and analyze the proposed reinforcement learning framework, keep optimizing the performance of the method, and apply it to solve more robotic challenges.

### v2

@@
-Vision-driven Autonomous Flight of UAV Along River Using Deep Reinforcement Learning with Dynamic Expert Guidance
+Synergistic Reinforcement and Imitation Learning for Vision-driven Autonomous Flight of UAV Along River
 Abstract
-Vision-driven autonomous flight and obstacle avoidance of Unmanned Aerial Vehicles (UAVs) along complex riverine environments for tasks like rescue and surveillance requires a robust control policy, which is yet difficult to obtain due to the shortage of trainable river environment simulators and reward sparsity in such environments.
-To easily verify the navigation controller performance for the river following task before real-world deployment, we developed a trainable photo-realistic dynamics-free riverine simulation environment using Unity.
-Successful river following trajectories in the environment are manually collected and Behavior Clone (BC) is used to train an Imitation Learning (IL) agent to mimic expert behavior and generate expert guidance.
-Finally, a framework is proposed to train a Deep Reinforcement Learning (DRL) agent using BC expert guidance and improve the expert policy online by sampling good demonstrations produced by the DRL to increase convergence rate and policy performance.
-This framework is able to solve the along-river autonomous navigation task and outperform baseline RL and IL methods.
-The code and trainable environments are available [^1] https://github.com/lijianwen1997/Deep-Reinforcement-Learning-with-Dynamic-Expert-Guidance.
+Vision-driven autonomous flight and obstacle avoidance of Unmanned Aerial Vehicles (UAVs) along complex riverine environments for tasks like rescue and surveillance requires a robust control policy, which is yet difficult to obtain due to the shortage of trainable riverine environment simulators.
+To easily verify the vision-based navigation controller performance for the river following task before real-world deployment, we developed a trainable photo-realistic dynamics-free riverine simulation environment using Unity.
+In this paper, we address the shortcomings that vanilla Reinforcement Learning (RL) algorithm encounters in learning a navigation policy within this partially observable, non-Markovian environment.
+We propose a synergistic approach that integrates RL and Imitation Learning (IL).
+Initially, an IL expert is trained on manually collected demonstrations, which then guides the RL policy training process.
+Concurrently, experiences generated by the RL agent are utilized to re-train the IL expert, enhancing its ability to generalize to unseen data.
+By leveraging the strengths of both RL and IL, this framework achieves a faster convergence rate and higher performance compared to pure RL, pure IL, and RL combined with static IL algorithms.
+The results validate the efficacy of the proposed method in terms of both task completion and efficiency.
+The code and trainable environments are available [^1] https://github.com/lijianwen1997/Synergistic-Reinforcement-and-Imitation-Learning.
 I INTRODUCTION
 Autonomous flight of UAVs over inland waterways has been investigated in applications like water sampling [1], inspection of irrigation system [2], and cooperative drone-boat navigation tasks [3, 4, 5].
-However, very few works addressed the along-river autonomous navigation tasks of UAVs, which can be applied in rescue, surveillance, and inspection missions over inland waterways.
-Due to seasonal changes of river channels’ geometry and water level, manually planned GPS waypoints may not suit for accurate river following missions.
-Besides, when navigating in cluttered riverine environments, UAVs also need to avoid obstacles like trees and bridges, while being above the water.
-To address the above challenges, a camera is used as the major perception source for the UAV to take advantage of information embedded in images and the recent development in image processing.
-To reduce the cost and safety issues for real-world experiments, we developed a series of riverine simulation environments using Unity [6] to verify the feasibility of controlling a UAV to fly along a river with only photo-realistic vision input.
-The impact of environmental forces on UAV is not considered in the simulation environment and the UAV action in the 3D space is discretized to alleviate the navigation challenge of operating without a map along rivers.
-Using the above simplification, we decouple the optimization of high-level local waypoint-based control policies or navigation planners for UAVs from mid-level velocity control, which is more suitable for the river following task due to its larger acceptable operation space compared to autonomous driving tasks for cars.
-We pre-train an Imitation Learning (IL) agent using human demonstrations to provide basic guidance, then combine Reinforcement Learning (RL) algorithm with an online IL algorithm that accumulates and learns from decent agent experiences, to boost the learning speed and final agent performance.
-This will overcome the challenge of obtaining a performant controller using only RL algorithms since the only positive reward is given when the UAV traverses any newly visited discretized river segment.
-This training framework is examined to outperform Proximal Policy Optimization (PPO), Behavior Clone (BC), and PPO with static BC guidance.
+However, few studies addressed the along-river autonomous navigation tasks of UAVs, which can be applied in rescue, surveillance, and inspection missions over inland waterways.
+Seasonal variations in river channel geometry and water levels render manually planned GPS waypoints less reliable for precise river following missions.
+Moreover, in cluttered riverine environments, UAVs must navigate while avoiding obstacles such as trees and bridges, all while maintaining a safe altitude above the water.
+A camera can be used as the major perception source for the UAV to take advantage of information embedded in images and the recent developments in image processing.
+To reduce the cost and safety issues for real-world experiments, we developed a riverine simulation environment using Unity [6] to verify the feasibility of controlling a UAV to fly along a river with only photo-realistic vision input.
+Figure 1: System architecture.
+A human expert collects good trajectories before training and the transitions are represented by blue arrows.
+$a^{H}_{t}$ denotes the human expert’s action.
+The transitions during training are represented by black arrows.
+$a^{E}_{t}$ denotes the IL expert’s action.
+River following can be classified as a track following task, which poses significant training challenges as the environment does not naturally conform to a Markov Decision Process (MDP) due to the agent’s partial observability and the non-Markovian reward function [7].
+Specifically, the first-person camera view of the UAV agent in the riverine scene may lack sufficient water pixels to guide the agent along the river, or it may contain bifurcated river channels, confusing the agent’s navigation decisions.
+Moreover, the UAV agent receives positive rewards only upon visiting new river segments, with no reward for idling above already visited segments.
+Consequently, the immediate reward depends on both past observations and actions, rather than solely on the current ones, rendering the task a Partially Observable Markov Decision Process (POMDP) with non-Markovian rewards.
+To address this challenge, combined approaches of RL and Learning from Demonstration (LfD) have been proposed [8, 9, 10].
+However, how to effectively leverage the complementary advantages of imitative learning from demonstrations and reinforcement learning from experience for the track following task remains an area requiring further research and exploration.
+To investigate into this problem, we pre-train an IL expert using human demonstrations to provide basic guidance, then combine a RL algorithm with the online-updating IL algorithm that accumulates and learns from decent agent experiences, to mimic the progressive expert behaviors while learning.
+In our experiments, Proximal Policy Optimization (PPO [11]) is chosen as the RL agent, Behavior Cloning (BC [12]) the expert.
+This training framework is examined to boost the learning speed and improve final agent performance in the track following tasks, outperforming PPO, StaticBC, DynamicBC and PPO with StaticBC guidance baselines.
 In summary, the main contributions of this paper are:
-  • The creation of the trainable photo-realistic riverine simulation environment, which bridges the sim2real gap for vision-driven autonomous navigation of UAVs in the river following task.
-  • The development of Deep Reinforcement Learning with a dynamic expert guidance framework that interactively trains Reinforcement Learning and Imitation Learning agents to reduce the training time and increase the overall performance.
+  • The development of synergistic reinforcement and imitation learning framework that interactively trains RL agent and IL expert to reduce the convergence time and increase the overall performance.
+  • The creation of the trainable photo-realistic riverine simulation environment, which is publicly available, and serves as a testbed for vision-driven autonomous navigation of UAVs in the river following task.
 II RELATED WORK
 Autonomous river following in a relatively simple curved river segment has been tested using a water segmentation mask from HSV threshed images [13, 14].
-This image processing technique is preferable in terms of its implementation simplicity but has vulnerabilities like unverified generalization ability to rivers under different geometrical and light conditions, as well as the high safety hazards and experimental costs.
-Game engines as simulation environments have brought more advantages for vision-driven Reinforcement Learning (RL) training, due to their flexibility in customizing environmental parameters, their low cost as a testbed before real-world deployment, and most importantly, their photo-realism to reduce the sim2real gap in air [9, 10], land [11, 12] (indoor/outdoor), and even underwater applications [13].
+This water mask extraction technique and the control method both require careful manual tuning, especially for complex riverine environments with obstacles and multiple river channels.
+To alleviate laborsome manual tuning of the river following controller, modern image embedding methods and RL algorithms can be investigated in a photo-realistic riverine simulator.
+Game engines as simulation environments have brought more advantages for vision-driven RL training, due to their flexibility in customizing environmental parameters, their low cost as a testbed before real-world deployment, and most importantly, their photo-realism to reduce the Sim2Real gap in air [15, 16], land [17, 18] (indoor/outdoor), and even underwater applications [19].
 To achieve a safer and more robust autonomous navigation controller in a diverse and adaptive riverine domain, researchers utilized Unreal Engine to train the agent for the river following task [20, 21].
-Their proof of the superiority of the Variational Auto Encoder (VAE) [16] for image encoding, over Linear Regression method or end-to-end controller training, is adopted in our work.
-However, the reset criteria and the demonstrative trajectory collection approaches of their river simulation environment are manual-based, the agent action is limited to 2D planes, and the environment is not publicly available.
-To overcome these difficulties, our Unity-based river simulator can be automatically trained and tested for the river following task, supports 3D action space and collision detection, and is made public as an OpenAI gym [17] environment, thanks to Unity ML-Agents Toolkit [6].
-Reinforcement learning(RL) and Imitation Learning(IL) are two powerful techniques with significant applications in the field of robotics.
-Robots can learn optimal behavior through trial and error through RL.
-Robots can learn to avoid obstacles [18], plan paths [19], and perform tasks like autonomous driving [20] through RL even in complicated, dynamic situations.
-On the other hand, IL [21] leverages expert demonstrations to train robots.
-By monitoring and imitating the actions of experts, robots can acquire complex skills rapidly.
-This is valuable in scenarios where human expertise is readily available, as it allows robots to learn efficiently without extensive programming [22].
-However, both of these approaches have some limitations.
-RL frequently requires a large number of interactions with the environment to learn optimal policies, which can be time-consuming and expensive.
-Recently, Researchers have combined RL and IL to accelerate the training of RL using IL’s guidance.
-For example, [23] proposed to use the imitative expert policy to guide the learning of the DRL agent by regularizing the divergence between the DRL agent’s policy and the expert policy.
-[24] proposed a framework that uses a loss function that combines behavior cloning and 1-step Q-learning losses to enable the transition from BC to RL without performance degradation.
-However, after the DRL agent has developed a policy that is similar to the expert policy if the DRL agent tries to update the policy to improve the performance, it will be constrained by the expert policy or the value penalty thus the improvement is hindered.
-On the other hand, as they have predominantly learned to imitate the demonstrations, robots trained through IL may not be able to adapt well to unexpected circumstances.
-DAgger [25] improves on BC by training on a dataset that better resembles the observations the expert policy is likely to encounter.
-However, the IL agent is also constrained by the expert.
-Thus to overcome these limitations, we propose an interactive framework called DRL with dynamic expert guidance that trains an RL agent using expert guidance and improves the expert policy online by sampling demonstrations produced by the DRL expert so that IL and RL can be improved jointly.
-(a) Comparison of image captured in Wildcat Creek, Indiana, USA (left) and image from Unity river environment (right) that are alike in overall components layout and texture appearance.
-(b) Orthographic images of the training (left) and testing (right) river simulation environments.
-Figure 1: Real and simulation river environments.
+Their proof of the superiority of the Variational Auto Encoder (VAE) [22] for image encoding, over end-to-end controller training, is adopted in our work.
+However, the reset criteria of their river simulation environment are manual-based, the agent action is limited to a 2D plane, and the environment is not publicly available.
+To overcome these difficulties, our Unity-based riverine simulator can be automatically trained and tested for the river following task, supports 3D action space and collision detection, and is made public as an OpenAI gym [23] environment, thanks to Unity ML-Agents Toolkit [6].
+RL and IL are potent techniques in robotics.
+RL enables robots to learn optimal behavior through trial and error, mastering tasks like obstacle avoidance [24], path planning [25], and autonomous driving [26] even in dynamic environments.
+Conversely, IL harnesses expert demonstrations to swiftly train robots, allowing them to mimic complex skills.
+This method is particularly valuable when human expertise is accessible, streamlining robot learning without extensive programming [27].
+However, both approaches have limitations.
+RL training can be time-consuming and costly, often demanding a substantial number of interactions with the environment to learn optimal policies.
+IL’s effectiveness is constrained by its reliance on static training datasets, making it challenging to generalize to novel scenarios.
+To address these issues, researchers have begun integrating RL and IL to enhance training efficiency and performance.
+For instance, DAgger [28] enhances BC by training on a dynamically updating dataset mirroring the observations the trained policy is expected to encounter.
+However, it necessitates continual online querying of optimal actions from a human expert.
+Other approaches include frameworks that use the guidance of IL to train RL [29, 30], regulating the discrepancy between the agent’s policy and the expert’s.
+In particular, [30] combines BC and 1-step Q-learning losses to smoothly transition from IL to RL without performance degradation.
+However, for these methods, the IL policy is static, and further improvements of RL agents are impeded by the constraints imposed by the expert policy.
+To overcome these limitations, we introduce an interactive framework called ”Synergistic Reinforcement and Imitation Learning.” This framework trains an RL agent using IL expert advice while simultaneously refining the expert policy through online sampling of high-quality demonstrations generated by the RL agent.
+By facilitating collaborative improvement between IL and RL methodologies, this approach enhances training efficiency and overall performance.
+(a) Comparison of an image captured in Wildcat Creek, Indiana, USA (left) and image from Unity river environment (right) that are alike in overall components layout and texture appearance.
+(b) Orthographic images of the riverine simulation environment and the CliffCircular environment.
+Figure 2: Real and simulated riverine environments, and CliffCircular environment (with two extra random cliff grids) for river/track following task.
 III SIMULATION
-III-A Photo-realistic River Environments
-The riverine simulation environment is built in Unity using River Auto Material 2019  [^2] https://assetstore.unity.com/packages/tools/terrain/r-a-m-2019-river-auto-material-2019-145937 and Mountain Environment  [^3] https://assetstore.unity.com/packages/3d/vegetation/mountain-environment-dynamic-nature-191851.
-The river spline tool is provided by the former package to support fast, flexible and scriptable river building with various realistic water textures and shaders.
-The latter package provides a huge library of scanned forest assets like trees, branches, logs, and rocks, which helps to enrich the biomes of the riverine scene, as well as makes it more alike the real-world river scene components and appearance, as shown in Figure 0(a).
+III-A Photo-realistic Riverine Environment
+The riverine simulation environment is built in Unity using River Auto Material 2019 and Mountain Environment .
+The river spline tool is provided by the former package to support fast, flexible, and scriptable river building with various realistic water textures and shaders.
+The latter package provides a huge library of scanned forest assets like trees, branches, logs, and rocks, which helps to enrich the biomes of the riverine scene, as well as makes it more alike the real-world riverine scene components and appearance, as shown in Fig.
+2(a).
 Aside from the realistic river and terrain trees, the riverine simulation environment features several additional aspects that not only mimic the general objects in real-world rivers, creeks, and streams, but also make the training of river following agents more difficult: 1) river channels with varied widths and depths, 2) river tributaries that diverge from the main water stream, 3) bridges over the river that can obstruct the navigable path of UAVs, and 4) sharp and slow turns.
-(Figure 2)
 Several environment design choices are made for better photo-realism and more robust training.
-Instead of continuous velocity control, discrete position control for UAV movement is used, which eliminates the efforts of controller dynamics adaptation in real-world riverine environments with various wind speeds, water surface effects, etc.
+Instead of continuous velocity control, discrete waypoint control for UAV movement is used, which eliminates the efforts of controller dynamics adaptation in real-world riverine environments with various wind speeds, water surface effects, etc.
 Thus we will name the UAV agent as the camera agent in the contents hereafter.
-The circular shape of rivers in our simulated environments is designed for better camera view coherence without seeing too many unrealistic un-modeled parts (sky, terrains, etc.) as the camera agent moves to any valid pose along the river.
-Besides, in the circular river environment, no matter where the camera agent is reset at the beginning of a training episode, the agent policy always has a chance to learn to traverse the whole river before ending this episode, making the collected trajectory lengths more uniform and independent of agent reset positions.
-(a) Varying widths and depths
-(b) Tributary
-(c) Bridge
-(d) Challenging U turns
-Figure 2: Key features of the Unity river simulation environment.
-TABLE I: Multi-discrete action values of the camera agent in the actual execution order, whose controlling logic is the same as remote controllers of common UAVs.
+The circular shape of the river in our simulated environment is designed for better camera view coherence without seeing too many unrealistic un-modeled parts (sky, terrains, etc.) as the camera agent moves to any valid pose along the river.
+The annulus design of the riverine environment guarantees consistent episodic rewards achievable regardless of the initial reset position of the camera agent at the beginning of each episode.
+TABLE I: Multi-discrete action values of the camera agent.
 Action 0 1 2
@@
 Leftward-rightward Translation (m) 0.5 0 -0.5
-III-B Camera Agent
+III-B Riverine Camera Agent
 The simulation agent, created using Unity ML-Agents Toolkit [6], is a camera that looks forward with a fixed pitch angle ($20$ degrees in our experiments) with respect to the drone frame.
-Since no dynamics is included in this simulation, and all control actions are discrete, the drone camera itself is viewed as the agent.
 The camera agent is controlled by a four-dimensional multi-discrete action that maps from the command inputs of the common remote controller of UAVs: 1) translation up and down - throttle, 2) horizontal rotation counter-clockwise and clockwise - yaw rate, 3) translation forward and backward - pitch, 4) translation leftward and rightward - roll.
 Each action branch has three discrete values $\{0,1,2\}$, where $1$ means no action on that branch.
-The actions are applied in the drone coordinate frame (instead of the camera frame) with the right-hand rule, where the positive X axis points along the drone head and the positive Z axis points upward to the sky.
 The specific values of movement caused by actions are in Table I.
-III-C Gym Environment
+III-C Riverine Gym Environment
 A river can be abstracted as a 3D Catmull-Rom spline [31] that is interpolated from several control points.
@@
 This prohibits the turnaround motion of the agent since the built river is circular so that the agent can finish the whole-river navigation without any turnaround.
-Even if the agent does not trigger the above two ”done” criteria, lingering around or doing continuous meaningless movements with no reward improvement (over 50 steps in our experiments) will also trigger environment reset, and so does collision with bridges.
+Lingering around or doing continuous meaningless movements with no reward improvement (over $50$ steps in our experiments) will also trigger environment reset, and so does collision with bridges.
 If the environment ”done” is triggered during training, the camera agent will be reset at a random point inside the spline-based bounding volume, with a random yaw near either spline tangent of the nearest spline point.
-The above criteria for the river following task can be better visualized in Figure 5.
-Figure 3: Demonstrative diagram of valid activity space and acceptable yaw range of camera agent in the river following task.
-$h_{1}=1m,h_{2}=15m,\alpha=150^{\circ}$ in our experiments.
-The drone picture is from  [^5] https://www.jaycar.com.au/r-c-fpv-drone-with-1080p-camera/p/GT4118.
+The above criteria for the river following task can be better visualized in Fig.
+3.
+Figure 3: Demonstrative diagram of valid activity space and acceptable yaw range of camera agent in the river following task. $h_{1}=1m,h_{2}=15m,\alpha=150^{\circ}$ in our experiments.
 The river central spline can be simplified as a sequence of connected line segments with nearly the same length.
@@
 The line segment where the camera agent’s current position projects onto is denoted as visited.
-If the camera agent has visited all the river splines, it will get the highest reward, then the environment will be reset.
+If the camera agent has visited all the river spline segments, it will get the highest reward, then the environment will be reset.
 For collision detection, a mesh collider is used for bridges, and a box collider with a side length of $0.5$m is used for the camera agent.
 The overall reward function for the autonomous river following task is given below.
-$r(s,a)=\begin{cases}-1,&\text{agent out of spline bounding volume}\\ -1,&\text{agent yaw diverts too much}\\ -1,&\text{agent collides with bridge}\\ -1,&\text{agent no progress over certain steps}\\ r_{shaped}&\text{agent covers un-visited spline segment}\\ 0,&\text{otherwise}\end{cases}$ (1)
-where $r_{shaped}$ is the percentage of newly-visited segments $n$ after the action $a$ over all segments number $N$, multiplied by a constant value, $10$ in this case.
-$r_{shaped}=10*n/N\$ (2)
-All negative reward conditions denote the reset of the environment (Equation 1).
-Besides, no final positive reward is given when the camera agent has visited all river spline segments since the amplified shaped reward, Equation 2, already gives informative guidance for the river following task.
-Due to the same reason, no environment reset is applied for backup movement or staying still of the agent.
+$r(s,a)=\begin{cases}-1,&\text{agent fails}\\ r_{shaped}&\text{agent covers un-visited spline segment}\\ 0,&\text{otherwise}\end{cases}$ (1)
+Agent fails when 1) out of spline bounding volume, 2) yaw diverts too much, 3) collides with bridge, and 4) no progress over certain steps.
+$r_{shaped}$ is the percentage of newly-visited segments $n$ after the action $a$ over all segments number $N$, multiplied by a constant value, $10$ in this case: $r_{shaped}=10*n/N$.
+All negative reward conditions denote the reset of the environment (Eq.
+1).
+III-D CliffCircular Environment
+CliffCircular (Fig.
+2(b)) is derived from the CliffWalking environment [32, 33] and tailored for the track following task.
+Like the camera agent in the riverine environment, the agent in CliffCircular must navigate along walkable grids without stepping onto cliffs until it circumnavigates the central cliff block.
+Additionally, CliffCircular presents partial observability and non-Markovian characteristics due to the local observation of cliffs surrounding the agent’s current grid, the flexibility in determining the circulation direction, and the nature of the track following task, which necessitates consideration of historical information.
+Figure 4: CliffCircular environment diagram. Central track grids, agent location and observation grids are laid on top of the color layers of obstacle and walkable grids.
+CliffCircular is a $12$x12121212 grid world, illustrated in Fig.
+4.
+The observation space is the flattened $5$x5555 grids around the agent indicating cliff.
+Action space includes five discrete operations: no operation, up, right, down and left.
+The $+1$ reward will be given if agent’s projection on the central track grid is newly visited, $-100$ if agent falls off cliff, otherwise 00.
+Episode will be reset if cliff is triggered or all central track grids are visited.
+At the beginning of each episode, the agent’s location and two additional cliff grids are randomly initialized.
+This setup encourages the policy to learn a safe and robust track following strategy.
+A TimeLimit environment wrapper of length $128$ is used.
 IV METHODOLOGY
-Figure 4: System architecture.
-A human expert collects good trajectory before the training and the transitions are represented by blue arrows.
-$a^{H}_{t}$ denotes the human expert’s action.
-The transitions during training are represented by black arrows.
-$a^{E}_{t}$ denotes the IL expert’s action.
-The proposed method is shown in Figure 4 and algorithm 1 to solve the river following task.
-It requires a demonstration dataset from a human expert, an IL agent to provide guidance, and an RL agent that explores the environment, updates the dataset and learns a control policy.
-To train an IL agent to learn from demonstration, a dataset $\mathcal{D}^{E}:\{\tau_{i}\}_{N}$ for $N$ successful trajectories is needed.
-Each trajectory is a sequence of state-action pairs $\tau={s_{1},a_{1},...s_{t},a_{t}}$.
+The proposed method, illustrated in Fig.
+1 and detailed in Algorithm 1 for solving the river following task, relies on several components: a demonstration dataset provided by a human expert, a BC expert offering guidance, and a PPO agent responsible for environment exploration, dataset updates, and policy learning.
+The BC expert is firstly trained on the demonstration dataset, $\mathcal{D}^{E}:\{\tau_{i}\}_{N}$ for $N$ successful trajectories, resulting in the final form of StaticBC and initial form of DynamicBC baselines.
+Each trajectory is a sequence of state-action pairs $\tau={s_{1},a_{1},...,s_{t},a_{t}}$.
 The goal of the expert policy $\pi^{E}$ is to maximize the log-likelihood over $\mathcal{D}^{E}$.
 […]
-To keep improving the expert policy’s performance during the DRL training, trajectories with episodic rewards larger than the reward threshold $T_{reward}$ will be appended to $\mathcal{D}^{E}$.
-$T_{reward}$ is $4$ in our experiments.
-Once the dataset is updated, the expert policy is retrained using the newest $2000$ steps from the dataset until it can mimic at least $95\%$ behaviors from the demonstration.
-Algorithm 1 DRL with Dynamic Expert Guidance
-Expert collects dataset $\mathcal{D}^{E}$
+To enhance the performance of the expert policy during PPO training, trajectories with episodic rewards exceeding the reward threshold $T_{reward}$ are added to dataset $\mathcal{D}^{E}$.
+$T_{reward}$ is set to $4$ in the riverine environment, $18$ in the CiffCircular environment.
+Evaluation of the PPO agent occurs every $10000$ training steps in the CliffCircular environment and every $5000$ steps in the riverine environment, with more frequent updates to the expert policy in the latter due to its higher complexity.
+In the simpler CliffCircular environment with multi-discrete observation, duplicate transitions are removed from dataset $\mathcal{D}^{E}$ to mitigate bias.
+The expert policy undergoes retraining using the latest $2000$ steps from the dataset for $20$ epochs.
+Algorithm 1 Synergistic Reinforcement and Imitation Learning
+Collect the transition dataset $\mathcal{D}^{E}$ by the human expert
 Train the IL policy $\pi^{E}$
-Initialize policy network $\phi$, value network $\psi$
+Initialize policy network $\phi$, value network $\psi$ of RL agent
 repeat
      Observe state $s_{t}$
-     Sample actions $a_{t}\sim\pi^{\phi}$ and $a_{t}^{E}\sim\pi^{E}$
+     Generate action $a_{t}$ using $\pi^{\phi}$
      Execute the action $a_{t}$ in the environment
@@
      Store transition $(s_{t},a_{t},s_{t+1},r_{t})$ in the agent buffer $\mathcal{D}^{R}$
-     if time to update DRL then
-         Sample a mini-batch $\mathcal{B}$
-         Calculate KL divergence $D_{KL}$
-         Calculate loss $\mathcal{L}$
-         Update value network $\psi$
-         Update policy network $\phi$
+     if time to update RL agent then
+         Sample a mini-batch $\mathcal{B}$ from $\mathcal{D}^{R}$
+         Generate actions $a_{t}^{E}$ using $\pi^{E}$
+         Calculate the cross entropy $H(\pi^{E},\pi^{\psi})$ (Eq. 4)
+         Calculate loss $\mathcal{L}$ (Eq. 5)
+         Update value network $\psi$ and policy network $\phi$
      end if
-     if time to evaluation DRL then
+     if time to evaluate RL agent then
          Sample trajectories using $\pi^{\phi}$
@@
               Append the trajectory $\tau$ in $\mathcal{D}^{E}$
-              Update the IL policy $\pi^{E}$
+              Remove repeated transitions from the $\mathcal{D}^{E}$
+              Update the IL policy $\pi^{E}$ (Eq. 2 )
          end if
@@
 until Convergence or maximum training steps
-The target of RL is to obtain a policy, defined as a distribution over actions conditioned on states $\pi(a_{t}|s_{t})$, to maximize the long-term discounted cumulative rewards
+The target of PPO is to obtain a policy, defined as a distribution over actions conditioned on states $\pi(a_{t}|s_{t})$, to maximize the long-term discounted cumulative rewards
 […]
 where $\tau$ is a trajectory and $p_{\pi}(\tau)$ is the distribution of the trajectory under policy $\pi$, and $T$ is the time horizon.
-Due to Proximal Policy Optimization(PPO)’s scalability and generalization ability, we use PPO [27] as a DRL algorithm to update the parameters of the policy network.
-In our framework, we use policy loss, value loss, and additional action loss to calculate the total loss and optimize the learning parameters.
-The policy loss constrains the biggest possible improvement step on a policy using the data we currently have.
-$\mathcal{L}_{policy}=\mathbb{E}\Bigl{[}\mathrm{min}{\Bigl{(}r_{t}(\theta)\hat{A_{t}},clip(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A_{t}}\Bigr{)}}\Bigr{]}$ (5)
-where $r_{t}(\theta)=\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{\theta}_{old}(a_{t}|s_{t})}$ denotes the probability ratio between the current and old policy.
-$\hat{A_{t}}$ is the advantage.
-The value function network $V_{\psi}$ is updated by the following loss function:
-$\mathcal{L}_{value}=\mathbb{E}\Bigl{[}\Bigl{(}R_{t}-V_{\phi}(s_{t})\Bigr{)}^{2}\Bigr{]}$ (6)
-where $R_{t}$ represents the returned value. To encourage the learned policy to be close to the policy learned from the demonstration, we can set a penalty term in the loss function
-$\mathcal{L}_{action}=D_{KL}(\pi^{\theta},\pi^{E})=\sum\pi^{\theta}(a_{t}|s_{t})\log\frac{\pi^{\theta}(a_{t}|s_{t})}{\pi^{E}(a_{t}|s_{t})}$ (7)
-where $D_{KL}$ is the KL divergence that measures the divergence between two distributions.
-Combining the above loss functions for the DRK with dynamic expert guidance becomes
-$\mathcal{L}=c_{policy}\mathcal{L}_{policy}+c_{value}\mathcal{L}_{value}+c_{action}\mathcal{L}_{action}$ (8)
+To encourage the PPO policy to be close to the policy learned from the decent demonstrations, we add a penalty term in the loss function
+$\mathcal{L}_{action}=H(\pi^{E},\pi^{\psi})=-\sum\pi^{E}(a_{t}|s_{t})\log\pi^{\psi}(a_{t}|s_{t})$ (4)
+where $H(\pi^{E},\pi^{\psi})$ is the cross entropy that measures the difference between the expert policy $\pi^{E}$ and the RL policy $\pi^{\psi}$.
+Combining the above loss functions, PPO+DynamicBC becomes
+$\mathcal{L}=w_{1}\mathcal{L}_{policy}+w_{2}\mathcal{L}_{value}+w_{3}\mathcal{L}_{action}$ (5)
+where we keep the policy loss $\mathcal{L}_{policy}$ and value loss $\mathcal{L}_{value}$ of PPO intact.
+$w_{1}$, $w_{2}$, and $w_{3}$ represent the weight constants associated with each loss.
+$w_{1}$ and $w_{2}$ are set to 1.
+$w_{3}$ is set to $1$ at the beginning.
+If the PPO achieves a mean score better than BC during evaluation, $w_{3}$ is set to $0.2$ to encourage exploration.
+PPO is updated every $1024$ training steps.
 V SIMULATION VERIFICATION
-One training and one testing river simulation environment 0(b) are created to train learning algorithms and test their generalization abilities.
-Specifically, these maps differ in the overall river length, variation of river widths and depths, number of (sharp) turns, bridges, tributaries, and sand islands, whereas the lighting conditions and terrain textures are unified to compare the performance of vision-driven controllers in learning geometrical patterns for the river following task.
-Observation images are resized to $128$ x $128$ for both encoding and RL training, and the encoded image is a vector of $1024$ float numbers.
-Variational Autoencoder [16] is obtained for each map by training 2000 randomly collected images from all maps up to the current training map’s difficulty level.
-This progressive update of the image encoder depends on the texture coherence across all maps, reduces the impact of the encoding module on RL training, and accounts for the downward compatibility of the trained agent.
+In the riverine environment, observation images are resized to $128$ x $128$ for both encoding and RL training, and the encoded image is a vector of $1024$ float numbers.
+VAE model is obtained by training $2000$ randomly collected images from the environment.
 The encoded mean vector (without Gaussian noise) is used as the observation input to RL algorithms.
-We ask an expert to execute the designed river following tasks in the simulator and collect actions and states as expert demonstration data.
-The UAV is randomly initialized in the valid activity space.
-The expert tries to correct the pose of the UAV and then controls the UAV to follow the river.
-Each trajectory is $50$ steps long.
-We collect $50$ trajectories in the medium environment as $\mathcal{D}^{E}$.
-The expert policy $\pi^{E}$ is trained by behavior clone using $\mathcal{D}^{E}$.
-After $50$ epochs of training, the expert policy $\pi^{E}$ has a success rate of $93.27\%$ to predict the expert action from the demonstration and is able to achieve a mean reward of $0.50$ in the training environment.
-The expert policy is used to generate reference actions during the training of RL agents.
-Training is done on a desktop computer with a GeForce 2080 GPU and an Intel Core i7-8700 CPU.
+$50$ trajectories each with $50$ steps are collected by a human expert to form the dataset $\mathcal{D}^{E}$.
+After $50$ epochs of training, the expert policy $\pi^{E}$ has a success rate of $93.27\%$ to predict the expert action from the demonstration and is able to achieve a mean reward of $0.60$ in the riverine environment.
+This expert policy will be used to generate reference actions during the training of RL agent.
+In the CliffCircular environment, $100$ trajectories are collected by an expert, which achieves a mean episode reward of approximately 00.
+The BC expert, pre-trained on this dataset, attains a reward of $-0.24$, indicating a suboptimal policy that is capable of accomplishing the track following objective but commits few errors occasionally.
 Code related to RL and IL is based on StableBaseline3 [34] and Imitation Learning Baseline Implementations [35].
@@
 VI-A Training Results
-The proposed methods together with baseline methods (BC, PPO, and PPO+StaticBC) are used to train a camera agent in the river following environment.
-All the methods have been trained for $14000$ steps for $2$ different seeds.
-Figure5 shows the training results of different methods in the training environment.
-The training curve is generated by averaging the training results from different seeds and represented by a solid line of the mean reward with ±$\pm 1$ standard error.
-The episodic reward is calculated by averaging $100$ past episodes and updated every $1024$ steps.
-The blue horizon line represents the average of $100$ episodic rewards of the expert policy, which is obtained from 50 episodes of evaluation in $2$ different seeds.
-PPO baseline slowly improves throughout the reinforcement learning phase but is still lower than the pre-trained BC baseline.
-The PPO+DaynamicBC initially returns rewards similar to PPO+StaticBC baseline.
+The proposed method (PPO+DynamicBC), together with baseline methods (StaticBC, DynamicBC, PPO, and PPO+StaticBC), are used to train the discrete-action agents in both environments.
+All the methods have been trained for $14000$ steps in the riverine environment and $10000$ steps in the CliffCircular environment for $5$ different seeds.
+Fig.
+LABEL:fig:training_result illustrates the training results of different methods in two environments.
+The training curve is derived by averaging results from $5$ different seeds and is depicted as a solid line representing the mean reward, with a shaded region around it indicating the standard error.
+The episodic reward is computed by averaging over $100$ past episodes, with updates occurring every $1024$ steps.
+Additionally, the green horizon line signifies the performance achieved by a pre-trained (BC) model.
+In the CliffCircular environment, the PPO baseline slowly improves and surpasses the pre-trained BC baseline after $85000$ steps.
+The PPO+DynamicBC initially returns rewards similar to PPO+StaticBC baseline.
+However, PPO+DynamicBC’s learning curve shows a higher increasing speed after $10000$ steps and finally reaches a mean episodic reward of $17.90$ at $85000$ steps, while the maximum possible reward is $20$.
+The best mean episodic reward of PPO+StaticBC is less than $11.31$.
+After $10000$ steps of training, PPO+DynamicBC’s mean episode length is $28.50$, while PPO+StaticBC and PPO’s mean episode lengths are $32.62$ and $88.40$, respectively.
+This suggests that the proposed method has acquired a more efficient policy for completing the task in fewer steps, whereas PPO+StaticBC has developed a less optimal policy, requiring an additional $4.12$ steps to complete the task.
+Evidently, PPO encounters challenges in learning an effective policy within this POMDP track following environment with non-Markovian rewards.
+In the river following environment, PPO baseline slowly improves throughout the reinforcement learning phase but is still lower than the pre-trained BC baseline.
+The PPO+DynamicBC initially returns rewards similar to PPO+StaticBC baseline.
 Both methods can get a reward at or above the BC after $25000$ steps.
-However, after the DRL agent’s performance surpasses the BC, PPO+DaynamicBC’s learning curve shows a higher increasing speed and finally reaches a mean episodic reward of $2.0$ at $105000$ steps, while the best mean episodic reward of PPO+StaticBC is less than $1.5$.
+However, after the PPO agent’s performance surpasses the BC, PPO+DynamicBC’s learning curve shows a higher increasing speed and finally reaches a mean episodic reward of $2.0$ at $105000$ steps, while the best mean episodic reward of PPO+StaticBC is less than $1.5$.
 PPO+StaticBC also shows multiple decrease trends during training, mainly because of the constraint of the static expert policy.
-PPO+DaynamicBC shows one decrease trend ($110000$ to $13000$) however it recovered to $1.7$ after $10000$ steps.
-On average, PPO+DaynamicBC’s score is $1.5$ higher than PPO, $0.6$ higher than the BC, and $0.4$ higher than the PPO+StaticBC.
-Figure 5: Training results of different approaches in the river following environment. The green horizontal line represents the performance of the pre-trained BC.
-VI-B Testing Results
-Figure 6: Episode reward distributions of four algorithms on training and testing environments.
+PPO+DynamicBC shows one decrease trend ($110000$ to $13000$) however it recovered to $1.7$ after $10000$ steps.
+On average, PPO+DynamicBC’s score is $1.5$ higher than PPO, $0.6$ higher than BC, and $0.4$ higher than PPO+StaticBC.
+Figure 5: Training results of different approaches in the CliffCircular environment and river following environment.
+VI-B Evaluation Results
+Figure 6: Episode reward distributions of four algorithms on CliffCircular-gym-v0 and riverine environments.
 The horizontal bar represents the mean, the vertical bar represents 25% and 75% percentiles, and the white dot is the median.
 Statistics from 50 episodes for each algorithm.
-Although both algorithms PPO+StaticBC and PPO+DynamicBC can achieve high episodic reward in the training map, the dynamic one still achieves higher mean and median episodic rewards in both maps and leads the static one in terms of the maximum episodic reward by a large margin in the testing map (Figure 6).
-This verifies the superiority of PPO+DynamicBC algorithm since the IL (BC) model is evolving itself while guiding RL (PPO) model, which acts as a better regularizer compared to the PPO+StaticBC algorithm.
-PPO baseline and BC algorithms perform the worst in both training and testing maps, due to shortage of runtime expert guidance and limited generalization ability, respectively.
-Figure 7: Trajectories of two algorithms PPO+StaticBC and PPO+DynamicBC starting from the same pose in the training map, in two experiments.
-Star symbols signify the agent position every 50 steps.
+As shown Fig.
+6, the reward and navigation efficiency are compared through violin plots.
+To generate the episodic reward distribution, all methods with the best weights are evaluated for $50$ epochs in the same seed.
+In the CliffCircular environment, BC has both high and low scores, resulting in a relatively low mean reward.
+PPO without guidance learns how to avoid collisions but cannot figure out the task objective.
+PPO+StaticBC is able to increase the chance of finishing the task and decrease the chance of collision compared to StaticBC.
+PPO+DynamicBC achieves the best result.
+It learns an optimal policy to finish the track following task while avoiding cliffs.
+In the riverine environment, BC and the vanilla PPO can never achieve any reward larger than $6$ while both algorithms PPO+StaticBC and PPO+DynamicBC can achieve higher episodic reward of $10$.
+Noticeably, PPO+DynamicBC achieves the highest mean and median episodic rewards compared to all other methods.
+This verifies the superiority of PPO+DynamicBC algorithm since the IL (BC) model is evolving itself while guiding RL (PPO) model, which acts as a better regularizer compared to PPO+StaticBC algorithm.
+Meanwhile, PPO and BC baseline algorithms have inferior performance due to the shortage of runtime expert guidance and limited generalization ability, respectively.
+The difference in training convergence and evaluation performance further illustrates the difficulty of the track following tasks.
+More importantly, this synergistic combination of PPO and BC is validated to largely boost the training speed and final model performance.
+Figure 7: Trajectories of two algorithms PPO+StaticBC and PPO+DynamicBC starting from the same pose in the Unity riverine environment, in two evaluation experiments.
+Star symbols signify the agent’s position every 50 steps.
 Trajectory color gets darker as the agent moves further away from the starting position.
-To qualitatively compare the navigation efficiency and failure modes of both dynamic and static algorithms, we plot their trajectories in the training map when starting at the same locations (Figure 7).
-The static algorithm agent can only navigate about $70$% of river length in both experiments, before it failed due to exceeding river spline volume (distracted by the tributary 0(b)) and bumping into the bridge, respectively.
+To qualitatively compare the navigation efficiency and failure modes of both dynamic and static algorithms, we plot their trajectories in the riverine environment when starting at the same locations (Fig.
+7).
+The static algorithm agent can only navigate about $70$% of river segments in both experiments, before it failed due to exceeding river spline volume (distracted by the tributary 2(b)) and bumping into the bridge, respectively.
 This is better visualized by the agent view near blue circles.
@@
 It also rejects tributary when seeing the main stream ahead by timely heading correction.
-Both cases can be viewed in figures near the red circles in Figure 7.
+Both cases can be viewed in figures near the red circles in Fig.
+7.
 VII CONCLUSION
-In this paper, we develop two riverine simulation environments to train a UAV agent to autonomously follow the river and avoid obstacles.
-A novel reinforcement learning method that utilizes the dynamic IL expert has been proposed.
-The performance of the proposed method has been compared with the baseline methods (PPO, BC, and PPO+StaticBC) in the riverine environments and achieves good qualitative and quantitative results.
-In the training environment, it achieves at least $0.4$ more rewards than all other methods on average during training and the highest best reward, and it is able to efficiently finish the river course without collision.
-It has good generalization ability as it also performs best in an unseen and difficult testing environment.
-In the near future, tunable environmental light conditions like light directions and intensities can be enabled in the simulated riverine gym environment to increase policy robustness in the way of domain randomization and transfer the learned river following ability to a real UAV platform and test it in real-world riverine environment.
-On the other hand, we will systematically study and analyze the proposed reinforcement learning framework, keep optimizing the performance of the method, and apply it to solve more robotic challenges.
+In this paper, we develop a photo-realistic riverine simulation environment to train a UAV agent to autonomously follow the river while avoiding obstacles.
+A novel framework that synergistically combines RL and IL is proposed, and evaluated in both a riverine environment and a customized grid world track following environment.
+The performance of the proposed method (PPO+DynamicBC), compared with the baseline methods (PPO, StaticBC, DynamicBC and PPO+StaticBC) in both environments, achieves the best qualitative and quantitative results.
+In the riverine environment, it achieves at least $0.4$ more rewards than all other methods on average, and the highest best and median rewards.
+In the CliffCircular environment, this algorithm finishes the track following task with nearly full score (episode reward $19.04$ out of $20$), with significantly lower reward variance than all baseline algorithms.
+Further research can improve the proposed method in several aspects.
+Firstly, instead of relying solely on the episodic reward threshold to filter trajectories, specific transitions can also be evaluated by some advantage measure and the valuable ones can be merged to the dataset for IL expert refinement.
+This could reduce the demonstration buffer size, bias of this dataset, and the overall training time.
+Secondly, IL expert’s policy might not consistently outperform that of the RL agent during training process.
+Therefore, when to utilize the action loss for updating the RL agent could be further investigated.
+Thirdly, observation augmentation (with historical observations or actions) techniques can be implemented to push the partially observable non-Markovian decision process to MDP for convergence guarantee in the track following tasks.